I have two probability questions that really through me a loop. Can someone answer the questions and show there work? I appreciate any help.
Thanks

1. Suppose you have 3 nickels, 2 dimes, and 6 quarters in your pocket. If you draw a coin randomly from your pocket, what is the probability that
a. You will draw a dime?
b. You will draw a half-dollar?
c. You will draw a quarter?

2. You are rolling a pair of dice, one red and one green. What is the probability of the following outcomes:
a. The sum of the two numbers you roll from the dice is 11.
b. The sum of the two numbers you roll is 7.
c. The sum of the two numbers you roll is 13.

Answer by longjonsilver(2297)

1.
nickel = n
dime = d
quarter = q
P(n) = 3 nickels out of 11 coins = 3/11
P(d) = 2 dimes out of 11 coins = 2/11
P(q) = 6 quarters out of 11 coins = 6/11
2.
Draw a little table of the 2 dice values and their sums (hope this comes out ok on here)...
___1_2_3__4__5__6
-----------------
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12
Looking at this, there are a total of 36 outcomes, so...
P(11) = 2/36 = 1/18
P(7) = 6/36 = 1/6
P(13) = 0
jon